# Geometric Multi-color Message Passing Graph Neural Networks for Blood-brain Barrier Permeability Prediction

### v2

@@ 1 Introduction
The blood-brain barrier (BBB) is a highly selective interface that regulates the exchange of substances between the bloodstream and the brain, ensuring central nervous system (CNS) homeostasis by preventing the entry of harmful compounds while allowing essential nutrients and gases to pass through [1, 2]. BBB permeability (BBBP) is tightly controlled by the endothelial cells that form the barrier, whose tight adhesion and adhesion junctions restrict the passage of large or hydrophilic molecules, allowing only small, lipophilic, or actively transported compounds to cross [2].
@@
Predicting BBB permeability is critical for CNS drug development. Compounds with insufficient permeability of BBB can fail to achieve therapeutic concentrations in the brain [3, 4, 5]. BBBP data can be broadly categorized into numerical and categorical types. Numerical data quantify a compound’s permeability by comparing its concentration in the brain to its concentration in the blood, providing a continuous scale for measuring how effectively a substance crosses the BBB. In contrast, categorical data offer a binary classification, labeling compounds as either BBB+ (permeable) or BBB- (non-permeable), which simplifies evaluation by indicating whether a substance can penetrate the barrier.
@@
Various in vivo and in vitro studies have been conducted to assess BBB permeability in drug molecules. In vivo, models provide physiologically relevant information but are resource-intensive, time-consuming, and ethically challenging [6]. In vitro models, including coculture and dynamic systems [7], offer controlled environments for the testing of compounds but still involve significant costs and time. To address these limitations, in silico methods have gained prominence as efficient alternatives for BBBP prediction. These models can be broadly categorized into two main types: traditional quantitative structure-activity relationship (QSAR) models and machine learning-based models.
@@
QSAR models are among the earliest computational approaches used to predict BBB permeability. These models establish statistical correlations between molecular descriptors (e.g., lipophilicity, molecular weight, and hydrogen bonding capacity) and experimentally determined BBB permeability. Early QSAR models relied on multiple linear regression (MLR) and partial least squares (PLS) to identify molecular features associated with BBB permeability [8, 9, 10]. Some QSAR models use physicochemical rules, such as Lipinski’s Rule of Five or Veber’s Rule[11, 12], to estimate whether a compound is likely to penetrate the BBB. While these QSAR models provide interpretability, they often struggle with complex, nonlinear relationships in molecular data, leading to reduced predictive accuracy for diverse chemical compounds [9].
@@
With the availability of large molecular datasets [13, 14], machine learning (ML) models have emerged as a more flexible alternative to QSAR-based methods. These models improve predictive performance by capturing non-linear relationships between molecular features and BBB permeability. Support vector machine (SVM), Random forest (RF), and gradient boosting trees (GBT) algorithms are particularly effective in handling high-dimensional molecular data, integrating a vector of molecular descriptors to enhance prediction accuracy [15, 12, 16, 17]. Deep learning architectures have further advanced this field by automating feature extraction. For instance, convolutional neural networks (CNNs) are applied to 2D molecular images, learning spatial patterns associated with BBB permeability [18]. Whereas recurrent neural networks (RNNs) and long short-term memory (LSTM) networks process SMILES representations to model molecular properties in a sequential manner [19]. Additionally, hybrid approaches such as the DeePred-BBB Model [16] and LightBBB framework [20] incorporate various descriptor-based features, including physicochemical properties and molecular fingerprints, to enhance classification performance.
@@
In recent years, graph neural networks (GNNs) have emerged as a promising alternative by shifting from descriptor-based input representations to direct processing of molecular graphs [21]. Unlike traditional ML/DL models that rely on precomputed molecular fingerprints or an array of descriptors, GNNs represent molecules as graphs, where atoms are treated as nodes and chemical bonds as edges, enabling the direct extraction of structural and relational features [22]. Message passing neural networks (MPNNs) [23] have been widely applied in this context, as these architectures allow information to be exchanged between atomic neighbors through iterative message-passing steps. GSL-MPP [24] extends standard MPNNs by integrating graph structure learning on a molecular similarity graph, utilizing both intra- and inter-molecular relationships. CoMPT [25] introduces a node-edge message interaction mechanism within a transformer-based framework to explicitly capture edge information. It also employs a message diffusion strategy to prevent over-smoothing and enhance long-range dependencies. CD-MVGNN [26] introduces a multi-view GNN that equally considers atoms and bonds by using dual encoders and a cross-dependent message-passing scheme for bidirectional information exchange.
@@
In this study, we introduce a novel geometric multi-color message-passing graph neural network (GMC-MPGNN) that extends the capabilities of standard MPNNs by incorporating atomic-level geometric features. Unlike conventional GNNs that operate purely on molecular connectivity, our proposed model integrates spatial representations and captures multi-range molecular interactions through the construction of multi-color subgraphs. Beyond classification tasks in BBBP, we extend the application of GMC-MPGNN to regression tasks using the B3DB dataset [13]. Our results demonstrate that GMC-MPGNN outperforms existing state-of-the-art deep learning models, setting a new benchmark for BBBP modeling.
@@ 2.1 Weighted Colored Subgraph Representation
In this work, we first represent molecules through weighted colored subgraphs (WCS) to capture the geometric and chemical properties critical for accurately predicting molecular interactions and permeability. To this end, we model the molecule of interest via a molecular graph $\mathcal{G(V,E)}$, where vertices $\mathcal{V}$ denote the atoms, colored by their atom types. Edges $\mathcal{E}$ represent the non-covalent interactions between atoms. For a given molecular dataset, we collect a set $\mathcal{C}$ containing all the distinct atom types, we denote $\mathcal{C}_{k}$ as a $k^{\text{th}}$ atom type in $\mathcal{C}$. In the current BBBP datasets, we consider 12 atom types, including C, H, O, N, P, Cl, F, Br, S, Si, I, and X, where X represents all other atom types. Thus, $_{1}=\text{C}$ C, $C_{2}=\text{H}$, and so on. Formally, vertices $\mathcal{V}$ can be expressed as:
@@
where $\mathbf{r}_{i}$ and $\alpha_{i}$ are the 3D coordinate and atom type for the $i^{\text{th}}$ atom, respectively. We define edge $\mathcal{E}$ using a weighted function $\Phi$, typically selected as a generalized exponential or Lorentz function [30], which captures the decay of interaction strength as interatomic distances increase:
=$\mathcal{E}=\left\{\Phi\left(\left\|\mathbf{r}_{i}-\mathbf{r}_{j}\right\|;\eta_{kk^{\prime}}\right)\mid\alpha_{k}=\mathcal{C}_{k},\alpha_{j}=\mathcal{C}_{k^{\prime}};i,j=1,2,\ldots,N\right\},$ (2)
where $\left\|\mathbf{r}_{i}-\mathbf{r}_{j}\right\|$ denotes the Euclidean distance between atoms $i$ and $j$, and $\eta_{kk^{\prime}}$ is the characteristic distance, defined as $\eta_{kk^{\prime}}=\eta\left(r_{\text{vdw}_{k}}+r_{\text{vdw}_{k^{\prime}}}\right)$, with $r_{\text{vdw}}$ representing the van der Waals radius and $\eta$ as a scalar parameter. Common choices for $\Phi$ include:
@@
where $\kappa$ is a parameter governing the decay rate of interaction strength, effectively approximating how quickly the interactions weaken with increasing distance.
We now construct a multi-color subgraph $\mathcal{G}_{kk^{\prime}}(\mathcal{V}_{kk^{\prime}},\mathcal{E}_{kk^{\prime}})\subset\mathcal{G(V,E)}$, for a pairwise atom types $\mathcal{C}_{k}$ and $\mathcal{C}_{k^{\prime}}$ with $\mathcal{C}_{k},\mathcal{C}_{k^{\prime}}\in\mathcal{C}$. Here vertices $\mathcal{V}_{kk^{\prime}}\subset\mathcal{V}$ corresponds to atom of types $\mathcal{C}_{k}$ and $\mathcal{C}_{k^{\prime}}$ in the original molecule and edges $\mathcal{E}_{kk^{\prime}}\subset\mathcal{E}$ account for the selected interactions between these atom types. We here consider two cases
@@
In all cases, we exclude covalent interactions where two atoms are too close to each other. To enforce this constraint, we set a cutoff such that if the distance between two atoms satisfies $\left\|r_{i}-r_{i}\right\|<r_{\text{vdw}_{i}}+$ $r_{\text{vdw}_{j}}+\sigma$, the interaction is ignored. Here, $\sigma$ is the standard deviation of all van der Waals radii in the given dataset.
The pairwise interaction of all the atoms in the subgraph $\mathcal{G}_{kk^{\prime}}$ can be represented in the weighted adjacency matrix as the following:
$\left(A_{kk^{\prime}}(\Phi,\eta_{kk^{\prime}})\right)_{ij}=\begin{cases}\Phi\left(\left\|\mathbf{r}_{i}-\mathbf{r}_{j}\right\|;\eta_{kk^{\prime}}\right),&\text{ if }(i,j)\in\mathcal{E}_{kk^{\prime}}\\ 0,&\text{ otherwise }\end{cases}$ (4)
The degree matrix $D_{kk^{\prime}}$ is a diagonal matrix where each entry represents the sum of the weights of the edges connected to a given node:
$\left(D_{kk^{\prime}}\right)_{ii}=\sum_{j\in\mathcal{V}_{kk^{\prime}}}\left(A_{kk^{\prime}}\right)_{ij}$ (5)
Finally, the Laplacian matrix $L_{kk^{\prime}}$ is computed as:
$L_{kk^{\prime}}=D_{kk^{\prime}}-A_{kk^{\prime}}$
The adjacency and Laplacian matrices capture the structural properties of the subgraph and are useful for further spectral and geometric graph analysis.
@@ 2.2 Weighted Colored Subgraph-based Atomic Features
To complement these subgraph-based descriptors, we integrate cheminformatics atomic features (CAF), denoted as $\mathcal{X}_{i}^{\text{CAF}}$, derived from cheminformatics tools such as RDKit [31]. These include the atom type, represented by a one-hot encoding of the atomic number (100 dimensions), the number of bonds the atom is involved in (6-dimensional one-hot encoding), formal charge (5-dimensional encoding), and chirality (4-dimensional encoding capturing cases such as unspecified, tetrahedral CW/CCW, or other). Additional descriptors include the number of bonded hydrogen atoms (5-dimensional), hybridization state (sp, sp2, sp3, sp3d, or sp3d2, encoded over 5 dimensions), aromaticity (a binary indicator of whether the atom is part of an aromatic system), and atomic mass (a real-valued scalar scaled by 1/100). All categorical features are one-hot encoded to ensure consistency and numerical stability in learning [32].
In addition to atomic-level descriptors, we incorporate bond features, $\mathcal{X}_{ij}^{\text{Bond}}$, between atoms $i$ and $j$ to enrich the message-passing process with edge-level information [32]. These features, also derived from RDKit, include the bond type (single, double, triple, or aromatic, encoded in 4 dimensions), whether the bond is conjugated (1-dimensional), whether the bond is part of a ring (1-dimensional), and the bond stereochemistry (none, any, E/Z or cis/trans, encoded in 6 dimensions). All bond features are represented using one-hot encoding.
@@ 3 Datasets Preparations
In this work, we utilize three distinct datasets related to BBBP. For clarity throughout the manuscript, we refer to them as $BBBP_{\text{cls}}^{\text{MolNet}}$ and $BBBP_{\text{cls}}^{\text{B3DB}}$ for the two classification datasets, and $BBBP_{\text{reg}}^{\text{B3DB}}$ for the regression dataset.
The dataset $BBBP_{\text{cls}}^{\text{MolNet}}$ is the MoleculeNet [14] classification dataset comprising a total of 2,039 compounds, with 1,560 classified as positive (BBB+) and 479 as negative (BBB–). We clean the SMILES strings by removing isolated ions such as [H+] or [Cl-], along with their associated prefixes. The 3D structures for this dataset are provided by WeiLab [33].
Both the second classification dataset, $BBBP_{\text{cls}}^{\text{B3DB}}$, and the regression dataset, $BBBP_{\text{reg}}^{\text{B3DB}}$, are derived from the B3DB benchmark introduced by Meng et al. [13]. The classification dataset initially contained 7,807 compounds, while the regression dataset included 1,051 compounds annotated with SMILES strings and logBB values.
Since neither dataset provides 3D molecular structures, we applied a consistent preprocessing pipeline to both. First, we cleaned the SMILES strings by removing isolated ions and then validated them using RDKit to ensure compatibility with graph-based models. Molecules that could not be parsed by RDKit were excluded 2 from the classification set and none from the regression set at this stage.
Next, to generate a single, low-energy 3D conformer for each molecule, we implemented a hierarchical pipeline. Our primary method utilized the OpenEye OMEGA Toolkit [34], a state-of-the-art tool that performs a high-quality, knowledge-based conformational search. For each input SMILES string, OMEGA was configured to generate a diverse ensemble of conformers and write the single, lowest-energy structure directly to a file in the required MOL2 format. For any molecules that failed to process with the OMEGA software, we employed a robust open-source fallback pipeline. This secondary procedure first used the RDKit library in Python to generate an initial 3D structure with the ETKDGv2 algorithm, which was saved as an intermediate PDB file. Subsequently, we used the OpenBabel [35] library to convert the PDB file into the final MOL2 format.
Despite this multi-tool approach, some conversions failed: 65 molecules in the classification set and 4 in the regression set could not be processed. As a result, we obtained 7,740 valid molecules for $BBBP_{\text{cls}}^{\text{B3DB}}$ and 1,047 for $BBBP_{\text{reg}}$, which were used in downstream experiments.
@@
To ensure a robust evaluation of our models, we employed a scaffold splitting method to partition each dataset into training, validation, and test sets. Scaffold splitting groups molecules by their core structures, helping to ensure structural diversity in the training set and providing a stringent test of generalizability on novel chemical scaffolds. We adopted an 8:1:1 ratio for all three datasets, as recommended in recent studies [36, 37]. This approach mimics real-world drug discovery scenarios, where models must predict the properties of previously unseen compounds.
For each dataset, we performed 21 scaffold-based splits using seed values from 0 to 20. Final model performance is reported as the average across all splits to ensure stability and robustness in our results.

### v4

@@ 1 Introduction
The blood-brain barrier (BBB) is a highly selective interface that regulates the exchange of substances between the bloodstream and the brain, ensuring central nervous system (CNS) homeostasis by preventing the entry of harmful compounds while allowing essential nutrients and gases to pass through [daneman2015blood, kadry2020blood]. BBB permeability (BBBP) is tightly controlled by the endothelial cells that form the barrier, whose tight adhesion and adhesion junctions restrict the passage of large or hydrophilic molecules, allowing only small, lipophilic, or actively transported compounds to cross [kadry2020blood].
@@
Predicting BBB permeability is critical for CNS drug development. Compounds with insufficient permeability of BBB can fail to achieve therapeutic concentrations in the brain [vilella2015endocytosis, gao2013targeted, dong2018current]. BBBP data can be broadly categorized into numerical and categorical types. Numerical data quantify a compound’s permeability by comparing its concentration in the brain to its concentration in the blood, providing a continuous scale for measuring how effectively a substance crosses the BBB. In contrast, categorical data offer a binary classification, labeling compounds as either BBB+ (permeable) or BBB- (non-permeable), which simplifies evaluation by indicating whether a substance can penetrate the barrier.
@@
Various in vivo and in vitro studies have been conducted to assess BBB permeability in drug molecules. In vivo, models provide physiologically relevant information but are resource-intensive, time-consuming, and ethically challenging [bagchi2019vitro]. In vitro models, including coculture and dynamic systems [cucullo2011dynamic], offer controlled environments for the testing of compounds but still involve significant costs and time. To address these limitations, in silico methods have gained prominence as efficient alternatives for BBBP prediction. These models can be broadly categorized into two main types: traditional quantitative structure-activity relationship (QSAR) models and machine learning-based models.
@@
QSAR models are among the earliest computational approaches used to predict BBB permeability. These models establish statistical correlations between molecular descriptors (e.g., lipophilicity, molecular weight, and hydrogen bonding capacity) and experimentally determined BBB permeability. Early QSAR models relied on multiple linear regression (MLR) and partial least squares (PLS) to identify molecular features associated with BBB permeability [patel2014quantitative, faramarzi2022development, platts2001correlation]. Some QSAR models use physicochemical rules, such as Lipinski’s Rule of Five or Veber’s Rule[lipinski2000drug, sakiyama2021prediction], to estimate whether a compound is likely to penetrate the BBB. While these QSAR models provide interpretability, they often struggle with complex, nonlinear relationships in molecular data, leading to reduced predictive accuracy for diverse chemical compounds [faramarzi2022development].
@@
With the availability of large molecular datasets [b3db, wu2017moleculenet], machine learning (ML) models have emerged as a more flexible alternative to QSAR-based methods. These models improve predictive performance by capturing non-linear relationships between molecular features and BBB permeability. Support vector machine (SVM), Random forest (RF), and gradient boosting trees (GBT) algorithms are particularly effective in handling high-dimensional molecular data, integrating a vector of molecular descriptors to enhance prediction accuracy [kumar2021b3pred, sakiyama2021prediction, kumar2022deepred, yuan2018improved]. Deep learning architectures have further advanced this field by automating feature extraction. For instance, convolutional neural networks (CNNs) are applied to 2D molecular images, learning spatial patterns associated with BBB permeability [tang2022merged]. Whereas recurrent neural networks (RNNs) and long short-term memory (LSTM) networks process SMILES representations to model molecular properties in a sequential manner [alsenan2020recurrent]. Additionally, hybrid approaches such as the DeePred-BBB Model [kumar2022deepred] and LightBBB framework [shaker2021lightbbb] incorporate various descriptor-based features, including physicochemical properties and molecular fingerprints, to enhance classification performance. Further advancements have also explored multi-task learning frameworks, which simultaneously predict several properties, and strategies involving the fusion of multi-level features to create more comprehensive molecular representations [yang2025multitask, yang2025multitasktoxicity].
@@
In recent years, graph neural networks (GNNs) have emerged as a promising alternative by shifting from descriptor-based input representations to direct processing of molecular graphs [wu2020comprehensive]. Unlike traditional ML/DL models that rely on precomputed molecular fingerprints or an array of descriptors, GNNs represent molecules as graphs, where atoms are treated as nodes and chemical bonds as edges, enabling the direct extraction of structural and relational features [kipf2017semisupervised]. Message passing neural networks (MPNNs) [gilmer2017neural] have been widely applied in this context, as these architectures allow information to be exchanged between atomic neighbors through iterative message-passing steps. GSL-MPP [gslmpp] extends standard MPNNs by integrating graph structure learning on a molecular similarity graph, utilizing both intra- and inter-molecular relationships. CoMPT [compt] introduces a node-edge message interaction mechanism within a transformer-based framework to explicitly capture edge information. It also employs a message diffusion strategy to prevent over-smoothing and enhance long-range dependencies. CD-MVGNN [cdmvgnn] introduces a multi-view GNN that equally considers atoms and bonds by using dual encoders and a cross-dependent message-passing scheme for bidirectional information exchange.
@@
In this study, we introduce a novel geometric multi-color message-passing graph neural network (GMC-MPGNN) that extends the capabilities of standard MPNNs by incorporating atomic-level geometric features. Unlike conventional GNNs that operate purely on molecular connectivity, our proposed model integrates spatial representations and captures multi-range molecular interactions through the construction of multi-color subgraphs. Beyond classification tasks in BBBP, we extend the application of GMC-MPGNN to regression tasks using the B3DB dataset [b3db]. Our results demonstrate that GMC-MPGNN outperforms existing state-of-the-art deep learning models, setting a new benchmark for BBBP modeling.
@@ 2.1 Weighted Colored Subgraph Representation
In this work, we first represent molecules through weighted colored subgraphs (WCS) to capture the geometric and chemical properties critical for accurately predicting molecular interactions and permeability. To this end, we model the molecule of interest via a molecular graph $\mathcal{G(V,E)}$, where vertices $\mathcal{V}$ denote the atoms, colored by their atom types. Edges $\mathcal{E}$ represent the non-covalent interactions between atoms. For a given molecular dataset, we collect a set $\mathcal{C}$ containing all the distinct atom types. We denote $\mathcal{C}_{k}$ as a $k^{\text{th}}$ atom type in $\mathcal{C}$. Formally, vertices $\mathcal{V}$ can be expressed as:
@@
where $\mathbf{r}_{i}$ and $\alpha_{i}$ are the 3D coordinate and atom type for the $i^{\text{th}}$ atom, respectively. We define edge $\mathcal{E}$ using a weighted function $\Phi$, typically selected as a generalized exponential or Lorentz function [nguyen2017rigidity], which captures the decay of interaction strength as interatomic distances increase:
$\mathcal{E}=\left\{\Phi\left(\left\|\mathbf{r}_{i}-\mathbf{r}_{j}\right\|;\eta_{kk^{\prime}}\right)\mid\alpha_{k}=\mathcal{C}_{k},\alpha_{j}=\mathcal{C}_{k^{\prime}};i,j=1,2,\ldots,N\right\},$ (2)
where $\left\|\mathbf{r}_{i}-\mathbf{r}_{j}\right\|$ denotes the Euclidean distance between atoms $i$ and $j$, and $\eta_{kk^{\prime}}$ is the characteristic distance, defined as $\eta_{kk^{\prime}}=\tau\left(r_{\text{vdw}_{k}}+r_{\text{vdw}_{k^{\prime}}}\right)$, with $r_{\text{vdw}}$ representing the van der Waals radius and $\tau$ is the scaling factor. This parameter allows the model to tune its focus on different interaction ranges. A value of $\tau<1$ prioritizes interactions at distances shorter than the sum of vdW radii, crucial for capturing strong, specific non-covalent forces like hydrogen bonds. Conversely, $\tau>1$ enables the model to learn from weaker, long-range electrostatic and van der Waals interactions. Common choices for $\Phi$ include:
@@
where $\kappa$ is the power parameter of the kernel, which helps approximate the ideal low-pass filter (ILF) [nguyen2017rigidity]. A high $\kappa$ value creates a sharp, steep decay, effectively filtering interactions to a very specific distance range. A low $\kappa$ value creates a “soft,” gradual decay, allowing a wider range of interactions to contribute. By optimizing $\kappa$, the model learns the most effective “shape” of the interaction potential.
We now construct a multi-color subgraph $\mathcal{G}_{kk^{\prime}}(\mathcal{V}_{kk^{\prime}},\mathcal{E}_{kk^{\prime}})\subset\mathcal{G(V,E)}$, for a pairwise atom types $\mathcal{C}_{k}$ and $\mathcal{C}_{k^{\prime}}$ with $\mathcal{C}_{k},\mathcal{C}_{k^{\prime}}\in\mathcal{C}$. Conceptually, this WCS approach creates separate subgraphs for specific types of interactions—for example, mapping how all nitrogen atoms interact with all oxygen atoms across the entire molecule. The strength (weight) of these interactions in $\mathcal{E}_{kk^{\prime}}$ depends on their spatial proximity and will allow the model to learn chemical properties that a simple connectivity graph would miss.
The vertices $\mathcal{V}_{kk^{\prime}}\subset\mathcal{V}$ correspond to all atoms of types $\mathcal{C}_{k}$ and $\mathcal{C}_{k^{\prime}}$ in the original molecule. We consider two cases for the edge set $\mathcal{E}_{kk^{\prime}}$:
@@
From these subgraphs, we compute intermediate matrices representing their structure. The detailed mathematical construction of the weighted adjacency ($A_{kk^{\prime}}$) and Laplacian ($L_{kk^{\prime}}$) matrices, along with the specific cutoff criteria for excluding covalent bonds, are provided in the Supporting Information (Section Methodological Details). Figure 1 illustrates this entire subgraph generation process.
@@ 2.2 Weighted Colored Subgraph-based Atomic Features
To complement these subgraph-based descriptors, we integrate cheminformatics atomic features (CAF), denoted as $\mathcal{X}_{i}^{\text{CAF}}$, derived from cheminformatics tools such as RDKit [landrum2006rdkit]. These include the atom type, represented by a one-hot encoding of the atomic number (100 dimensions), the number of bonds the atom is involved in (6-dimensional one-hot encoding), formal charge (5-dimensional encoding), and chirality (4-dimensional encoding capturing cases such as unspecified, tetrahedral CW/CCW, or other). Additional descriptors include the number of bonded hydrogen atoms (5-dimensional), hybridization state (sp, sp2, sp3, sp3d, or sp3d2, encoded over 5 dimensions), aromaticity (a binary indicator of whether the atom is part of an aromatic system), and atomic mass (a real-valued scalar scaled by 1/100). All categorical features are one-hot encoded to ensure consistency and numerical stability in learning [heid2023chemprop].
In addition to atomic-level descriptors, we incorporate bond features, $\mathcal{X}_{ij}^{\text{Bond}}$, between atoms $i$ and $j$ to enrich the message-passing process with edge-level information [heid2023chemprop]. These features, also derived from RDKit, include the bond type (single, double, triple, or aromatic, encoded in 4 dimensions), whether the bond is conjugated (1-dimensional), whether the bond is part of a ring (1-dimensional), and the bond stereochemistry (none, any, E/Z or cis/trans, encoded in 6 dimensions). All bond features are represented using one-hot encoding.
@@ 3 Datasets Preparations
In this work, we utilize three distinct datasets related to BBBP. For clarity throughout the manuscript, we refer to them as $BBBP_{\text{cls}}^{\text{MolNet}}$ and $BBBP_{\text{cls}}^{\text{B3DB}}$ for the two classification datasets, and $BBBP_{\text{reg}}^{\text{B3DB}}$ for the regression dataset. For all datasets, we first performed a cleaning step on the SMILES strings to remove isolated ions and ensure chemical validity.
The $BBBP_{\text{cls}}^{\text{MolNet}}$ dataset contains 2,039 compounds with pre-existing 3D structures. The larger $BBBP_{\text{cls}}^{\text{B3DB}}$ and $BBBP_{\text{reg}}^{\text{B3DB}}$ datasets, derived from the B3DB benchmark, did not include 3D coordinates. For these, we generated a single, low-energy 3D conformer for each molecule using a hierarchical pipeline. This process was primarily conducted with the OpenEye OMEGA Toolkit, with a fallback procedure using RDKit and OpenBabel to maximize coverage. After removing molecules that could not be processed, we obtained final datasets of 7,740 molecules for classification and 1,047 for regression. A full description of the dataset preparation is available in the Supplemental Information [SI_GMC_MPNN]
@@
To ensure a robust evaluation of our models, we employed a scaffold splitting method to partition each dataset into training, validation, and test sets. Scaffold splitting groups molecules by their core structures, helping to ensure structural diversity in the training set and providing a stringent test of generalizability on novel chemical scaffolds. We adopted an 8:1:1 ratio for all three datasets, as recommended in recent studies [shen2023molecular, chan2019advancing]. This approach mimics real-world drug discovery scenarios, where models must predict the properties of previously unseen compounds.
For each dataset, we performed 5 scaffold-based splits using seed values from 0 to 4. Final model performance is reported as the average across all splits to ensure stability and robustness in our results.